OpenStudy (anonymous):

You put $5000 in a savings account which earns compound interest at a rate of 0.5% per month. If you don't make any deposits or withdrawls, how much will you have in your account in 30 years (360 months)? Round your answer to the nearest whole cent. A. $21,609.71 B. $5807.00 C. $30,112.88 D. $22,324.85

OpenStudy (aum):

Here is the compound interest formula: \(\Large A = P(1 + \frac{ r }{ n })^{nt}\) A = Amount at maturity P = Principal Amount = $5,000 r = Annual interest rate in decimal = 12*.5 = 6%. In decimal: 0.06 n = compounding period (compounded how many times a year) = 12 t = years invested = 30 \(\large A = 5,000(1 + \frac{ 0.06 }{ 12 })^{12*30} = 5,000(1.005)^{360} = $30,112.88\)

OpenStudy (anonymous):

thank you:)

OpenStudy (aum):

You are welcome.